A machine gun fires 1200 bullet of mass 5g each in one min with speed 400m\s . Force required to hold gun is?
 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Force required is = 1200 *5 * 400/ ( 60 * 1000) ) = 40 n

Regards
Vikas TU
4 years ago
Dear student 
Explanation
Acceleration = v/t 
v = 400m/sec 
a = 400 /60 
Force required = m*a 
Total no. of Bullets = 1200 
Force required = 1200* 5/1000 * 400/60 
=> 40N 
H